# Forecasting Company Fundamentals

### v1

@@ Abstract
Company fundamentals are key to assessing companies’ financial and overall success and stability. Forecasting them is important in multiple fields, including investing and econometrics. While statistical and contemporary machine learning methods have been applied to many time series tasks, there is a lack of comparison of these approaches on this particularly challenging data regime. To this end, we try to bridge this gap and thoroughly evaluate the theoretical properties and practical performance of 22 deterministic and probabilistic company fundamentals forecasting models on real company data. We observe that deep learning models provide superior forcasting performance to classical models, in particular when considering uncertainty estimation. To validate the findings, we compare them to human analyst expectations and find that their accuracy is comparable to the automatic forecasts. We further show how these high-quality forecasts can benefit automated stock allocation. We close by presenting possible ways of integrating domain experts to further improve performance and increase reliability.
@@ 1 Introduction
Specifically for a value or factor investor, knowing future company fundamentals would be enormously beneficial in making successful investment decisions measured in total share price appreciation. For example, it has been shown that if these CFs had been known prospectively, investments could have been made much more successfully (Alberg & Lipton, 2018; Chauhan et al., 2020; Downey, 2020). For instance, using a clairvoyant model that always “predicts” the correct value without error, Chauhan et al. are able to reach a simulated annualized return of $\sim 40$% (for one year of looked-ahead). This naturally leads to the question of whether reliably forecasting CFs using machine learning models can enhance investment decisions. This is particularly promising since the intersection of machine learning (ML) and quantitative financial analysis has opened new possibilities for predicting company fundamentals, a domain traditionally dominated by statistical methods and manual assessments by analysts with very specific domain expertise. When instead of the clairvoyant model Chauhan et al. (2020) used real forecasts with inevitable prediction errors, this improved their lookahead factor model’s annualized return by 3.7 percentage points over a quantitative factor model based on merely current CFs. On the other hand, the work by Downey (2020) did not show a significant benefit of using automated forecasts. This uncertain yet promising state of affairs motivates forecasting CFs as precisely and reliably as possible to serve as building blocks for factor models. Subsequently, we can rigorously assess the benefit of these prognoses to factor investing as a case study.
@@
    An investigation of the model’s fitness for forecasting company fundamentals, a comparison to human forecasts, the availability of uncertainty estimates, their applicability to investment strategies by factor model backtesting.
@@ 2 Related Work
There have been some prior attempts at forecasting fundamentals. Most notably, experiments with an idealized clairvoyant model always predicting the known future and, subsequently, imperfect LSTMs showed that lookahead factor models (LFMs) based on CF forecasts can improve returns of portfolios (Alberg & Lipton, 2018; Chauhan et al., 2020). Sun (2019) studied fundamentals forecasts with a few simple deep learning (DL) models yet did not apply them to stock selection. Following the work of Alberg & Lipton (2018), Downey (2020) first replicated the “crystal ball” portfolio based on perfect fundamental forecasts and confirmed its impressive hypothetical performance for stock allocation. However, none of the investigated models (Random Forests, Gradient Boosting, Support Vector Machines, Multilayer Perceptron, Linear Regression) could forecast CFs with sufficient reliability to affect investing positively. Guerard et al. (2015) used analysts’ expectations on companies’ revenue and found this to be a strongly beneficial signal for stock selection. This again highlights the need for accurate company fundamental forecasts.
@@ 3 Model Selection for Forecasting
Table 1: Overview of the models considered in this evaluation. It shows if they are local, global, or global deep learning (DL) models. Further properties are if they are autoregressive (AR); the number of predicted variates; their use of additional covariates, static variables, or RevIN; and their ability to provide probabilistic forecasts.
@@
In addition to the preprocessing steps described later in Section 4.1.2, we evaluated applying Reversible Instance Norm (RevIN) (Kim et al., 2022) in all non-recurrent deep learning models as shown in Table 1. It removes some nonstationarity and possible shifts in distribution. RevIN removes the mean and rescaling to unit variance along each individual time series and variate $x$. Furthermore, the result is then transformed by an affine transform with learnable $\gamma$ and $\beta$: $\operatorname{RevIN}(x_{t})=\gamma\left(\frac{x_{t}-\mathbb{E}\left[x\right]}{\sqrt{\mathrm{Var}\left[x\right]}+\epsilon}\right)+\beta.$ We found it to be mostly beneficial, as shown quantitatively in detail in Section A.4, and will thus apply it throughout the paper.
@@ 3.1 Local Models
This simple baseline model repeatedly predicts the mean value of each variate it has seen during training: $\bm{x}_{T+h}=\frac{1}{T}\sum_{i=1}^{T}\bm{x}_{i}$. It is included as the arguably most simple model one should consider.
@@
In practice, it is often unclear how $p$, $d$, and $q$ shall be chosen. AutoARIMA introduces an automated algorithm for automatically determining them in a data-driven search (Hyndman & Khandakar, 2008).
@@ 3.2 Global Models
While not being the first ever RNNs, Long Short-term Memory models (LSTMs) (Hochreiter & Schmidhuber, 1997) popularized the concept of iteratively consuming an input sequence to produce both an output and a hidden state propagated to the next step. Using several gating functions, very deep LSTM networks can be trained via backpropagation through time. Gated recurrent units (GRUs) simplify LSTMs, often reaching similar performance with fewer parameters (Cho et al., 2014).
@@
It is often beneficial to decompose challenging tasks into a sequence of easier ones. N-BEATS, therefore, comprises a stack of doubly residually connected blocks, where each block performs both a forecast and a backcast (Oreshkin et al., 2019). The backcast is subtracted from the input and only then passed to the subsequent block, in effect removing the component used by the block and only leaving the following block with the residual. Similarly, the complete forecast is computed as the sum of individual forecasts, effectively modeling it as an additive decomposition. Each back- and forecast is computed based on predicting basis function coefficients to ease training and allow for inductive biases. We use the multivariate extension of Darts (Herzen et al., 2022).
@@ 4.1.1 Selected Indicators
We experimented with providing the model with additional static context for each company. Namely, we added region information (Americas/Europe/Japan/Rest of Asia) and sector classifications according to the Global Industry Classification Standard (GICS) by MSCI and S&P. We encoded them as numeric one-hot statics for the global models that can use it (see Table 1). However, we only retained the latter since the region information did not noticeably change the forecasts. This is likely due to it being very coarse and of limited relevance in globally diversified companies.
@@ 4.1.2 Data Preparation
The primary quarterly company fundamentals data was obtained by integrating several commercial data sources from S&P Global. Considerable manual preprocessing was required to bring information from balance sheets, income statements, and cash flow statements into a single format. In particular, restatements and pro forma/hypothetical statements introduced significant inconsistencies and jumps in the data if not accounted for appropriately. All values were converted into Euros for the sake of comparability. Companies were included if (1) they were publicly traded throughout the entire period and, therefore, published reports, and (2) they had at least 1 billion Euros in market capitalization once in the time span. Smaller businesses were excluded because they typically show different dynamics than huge ones. An additional 19 companies were removed due to highly atypical behavior. This left 2527 companies in the period from 2009 Q1 to 2023 Q3.
Appropriate data representation and normalization are critical in machine learning applications. In the particular case of CFs, one has to counteract the different orders of magnitudes of the indicators to ensure more similar numerical ranges of the data. We, therefore, propose an effective domain-specific normalization. For each company individually, all variates derived from the income statement are normalized by Total Revenues, and all variates from the balance sheet are normalized by Total Assets. Total Revenues and Total Assets are divided by each other, respectively. This proved to be more effective than a generic Yeo–Johnson power transformation (Yeo & Johnson, 2000). Finally, the data was normalized to have zero mean and unit variance per feature over all time steps and companies. Details on the resulting challenging dataset, including its moments, high degree of non-stationarity, and diverse seasonality, are provided in Section A.2.
@@ 4.1.3 Results
Figure 1: The error of different deterministic models when forecasting the normalized features. Lower is better. It shows the mean and standard deviation over all five target features, forecast horizon steps, and folds.
@@
The comparison of the models is shown visually in Figure 1 and in full detail in Table 2. In both metrics, the models fare notably differently. Overall, global models tend to give better results in terms of MSE, while the picture is more varied if performance is measured in sMAPE. Specifically, the univariate local models $\operatorname{ARMA}(4,4)$ and Prophet fail to provide consistent forecasts, as their high MSE and considerable variance hint at some major mispredictions. While in terms of MSE, the two deep learning RNN models (LSTM and GRU) give the best results, the simplistic local model $\operatorname{ARMean}(1)$— somewhat surprisingly—is best in terms of sMAPE. This highlights the need for forecasts that are not only accurate but reliable, too. Therefore, we will now go over to probabilistic forecasts.
Table 2: Comparison of average performance and standard deviation of the models measured by different metrics. The average and standard deviation is taken over all five target features, forecast horizon steps, and folds. The best is marked as bold and the runner-ups as underlined.
@@
To assess the reliability of the forecasters, we evaluated their ability to quantify uncertainty correctly. Not all models can estimate uncertainty, as shown in Table 1. Sampling forecasts can be achieved in different ways. For instance, one can introduce noise in the state space as in the $\operatorname{ARMA}$, $\operatorname{ARIMA}$, and AutoARIMA models or in the trend component of the Prophet model, which is similar to Monte Carlo dropout in deep learning models (Gal & Ghahramani, 2016). The remaining models can directly estimate the parameters of probability distributions. Therefore, and for a fair comparison, all remaining models perform quantile regression, where the boundaries of certain quantiles were regressed on. Note that these models are trained separately from the ones in the deterministic setting, even though represented jointly in Table 2. We have drawn 100 samples from each probabilistic model and evaluated the resulting empirical distribution to the known test data using the continuous ranked probability score (CRPS) (Gneiting & Raftery, 2007). Intuitively, we first define a helper cumulative distribution function $\mathbf{1}_{\{x\geq y\}}$ over $x$ that steps from zero to one at the scalar value $y$ that we consider the ground truth. The CRPS then measures the integrated squared difference between this helper and the true distribution $E(x)$. It is defined as
@@
Figure 2: The error of different models when forecasting the normalized features probabilistically. Lower is better. It shows the mean and standard deviation over all five target features, forecast horizon steps, and folds. Since the nCRPS score reduces to the absolute error (MAE) in the case of deterministic forecasts (i.e., a single sample), we include the non-deterministic models for direct comparison, marked with .
An example forecast transformed back to the original feature domain is provided in Figure 3. The normalized forecasts were inverse-transformed back into the original domain using the last-know training data to avoid any information leakage. The model successfully captures the overall dynamics of the variates while tending to lag slightly behind. In particular, the variance of the model appropriately captures the variability of the ground-truth time series.
@@
As shown in Figure 4, not all features are equally predictable by all models. For instance, the Total Equity is easily forecasted by $\operatorname{ARMean}(1)$, yet much harder for other models. This is plausible, given that it only moderately changes over time and thus might cause more overfitting. On the other hand, it is surprising that Total Revenues is relatively straightforward to predict for the deep learning models, indicating that some other covariates are predictive for it. The fact that the Operating Income is more challenging to predict suggests that the cost of revenues varies significantly, affecting the variate in ways that are hard to predict. Consequently, the Net Income is affected, albeit reduced, possibly due to more steady constant taxation and interests. By cumulating it over time, one derives the therefore equally predictable Cash from Operations.
Figure 4: Models differ significantly in the features they can successfully forecast. This plot shows the mean nCRPS/MAE per model and feature over all folds. Lower is better. It is apparent that the Operating Income and the Total Equity pose the greatest challenges to the models under investigation.
Global Black Swan events like the COVID-19 pandemic starkly impact forecasting performance, as shown in Figure 5. Note that for most variates and models this extends far beyond the difficulties in forecasting in the respective periods, like here, 2020 Q1 and onward. On the one hand, starting with 2019 Q2, the forecast evaluation is already infected with COVID-19 effects since we forecast one year into the future. Furthermore, the abnormal data from the pandemic period is then used to learn forecasts for later times, which no longer share the same pandemic dynamics, extending the effect beyond the end of the special period. We can also see that models excel in different parts of the time axis. For example, while $\operatorname{ARMean}(1)$ appeared to be overall decent in Figure 2, it produces terrible forecasts in the more recent time spectrum when we consider uncertainty estimation for Net Income while maintaining excellent performance on Total Euqity. There seems to be a visibly recurring pattern in predictability in Cash from Operations every four quarters, where the last quarter of each year is the least predictable.
@@
Figure 6: This shows the general trend of increasing forecast error with increasing look-ahead horizon, with a noticeable drop at four steps. The error is over all features and folds, where lower is better. The shaded area indicates one standard deviation. See Figure 5 for the legend.
@@
We also compared the machine forecasts with the ones of human expert analysts. To this end, we used data from StarMine by Refinitiv, which contains one-year-ahead non-probabilistic forecasts of the expected Total Revenues for 87.96% of the companies from 2013 Q3 to 2022 Q2. Thus, we need to compare them only with the automatic forecasts for that feature, those companies, the last of the four look-ahead steps, and the valid time window. The results are provided in complete form in Section A.5. In summary, automatic forecasts are superior for most metrics, even though different models than in Table 2 are leading. However, human analysts make mispredictions that better match the scale of the data, e.g., minor errors on small scales and large ones on bigger ground-truth numbers. This results in much better MAPE and sMAPE scores and might hint at potentially more calibrated uncertainty estimates if they were available. Given that human forecasts are commonly employed in investing (Guerard et al., 2015), it is interesting to see how their utility compares to automatic forecasts.

### v2

@@ Abstract
Company fundamentals are key to assessing companies’ financial and overall success and stability. Forecasting them is important in multiple fields, including investing and econometrics. While statistical and contemporary machine learning methods have been applied to many time series tasks, there is a lack of comparison of these approaches on this particularly challenging data regime. To this end, we try to bridge this gap and thoroughly evaluate the theoretical properties and practical performance of 24 deterministic and probabilistic company fundamentals forecasting models on real company data. We observe that deep learning models provide superior forecasting performance to classical models, in particular when considering uncertainty estimation. To validate the findings, we compare them to human analyst expectations and find that their accuracy is comparable to the automatic forecasts. We further show how these high-quality forecasts can benefit automated stock allocation. We close by presenting possible ways of integrating domain experts to further improve performance and increase reliability.
@@ 1 Introduction
Specifically for a value or factor investor, knowing future company fundamentals would be enormously beneficial in making successful investment decisions measured in total share price appreciation. For example, it has been shown that if these CFs had been known prospectively, investments could have been made much more successfully (Alberg & Lipton, 2018; Chauhan et al., 2020; Downey, 2020). For instance, using a clairvoyant model that always “predicts” the correct value without error, Chauhan et al. are able to reach a simulated annualized return of $\sim 40$% (for one year of looked-ahead). This naturally leads to the question of whether reliably forecasting CFs using machine learning models can enhance investment decisions. This is particularly promising since the intersection of machine learning (ML) and quantitative financial analysis has opened new possibilities for predicting company fundamentals, a domain traditionally dominated by statistical methods and manual assessments by analysts with very specific domain expertise. When instead of the clairvoyant model, Chauhan et al. (2020) used real forecasts with inevitable prediction errors, this improved their lookahead factor model’s annualized return by 3.7 percentage points over a quantitative factor model based on merely current CFs. On the other hand, the work by Downey (2020) did not show a significant benefit of using automated forecasts. This uncertain yet promising state of affairs motivates forecasting CFs as precisely and reliably as possible to serve as building blocks for factor models. Subsequently, we can rigorously assess the benefit of these prognoses to factor investing as a case study.
@@
    An investigation of the model’s fitness for forecasting company fundamentals, a comparison to human forecasts, the availability of uncertainty estimates, and their applicability to investment strategies by factor model backtesting.
@@ 2 Related Work
There have been some prior attempts at forecasting fundamentals. Most notably, experiments with an idealized clairvoyant model always predicting the known future and, subsequently, imperfect LSTMs showed that lookahead factor models (LFMs) based on CF forecasts can improve returns of portfolios (Alberg & Lipton, 2018; Chauhan et al., 2020). Sun (2019) studied fundamentals forecasts with a few simple deep learning (DL) models, yet did not apply them to stock selection. Following the work of Alberg & Lipton (2018), Downey (2020) first replicated the “crystal ball” portfolio based on perfect fundamental forecasts and confirmed its impressive hypothetical performance for stock allocation. However, none of the investigated models (Random Forests, Gradient Boosting, Support Vector Machines, Multilayer Perceptron, Linear Regression) could forecast CFs with sufficient reliability to affect investing positively. Guerard et al. (2015) used analysts’ expectations on companies’ revenue and found this to be a strongly beneficial signal for stock selection. This again highlights the need for accurate company fundamental forecasts.
@@ 3 Model Selection for Forecasting
Table 1: Overview of the models considered in this evaluation. It shows if they are local, global, global deep learning (DL), or global pretrained (PT) models. Further properties are if they are autoregressive (AR); the number of predicted variates; their use of additional covariates, static variables, or RevIN; and their ability to provide probabilistic forecasts.
@@
xLSTM-Mixer Kraus et al. (2024a) DL no multi
Chronos Ansari et al. (2024) PT yes uni
In addition to the preprocessing steps described later in Section 4.1.2, we evaluated applying Reversible Instance Norm (RevIN) (Kim et al., 2022) in all non-recurrent deep learning models as shown in Table 1. It removes some nonstationarity and possible shifts in distribution. RevIN removes the mean and rescales to unit variance along each individual time series and variate $x$. Furthermore, the result is then transformed by an affine transform with learnable $\gamma$ and $\beta$:
$\operatorname{RevIN}(x_{t})=\gamma\left(\frac{x_{t}-\mathbb{E}\left[x\right]}{\sqrt{\mathrm{Var}\left[x\right]}+\epsilon}\right)+\beta.$
We found it to be mostly beneficial, as shown quantitatively in detail in Section A.4, and will thus apply it throughout the paper.
@@ 3.1 Local Models
This simple baseline model repeatedly predicts the mean value of each variate it has seen during training: $\bm{x}_{T+h}=\frac{1}{T}\sum_{i=1}^{T}\bm{x}_{i}$. It is included as the arguably simplest model one should consider.
@@
In practice, it is often unclear how $p$, $d$, and $q$ shall be chosen. AutoARIMA introduces an automated algorithm for determining them in a data-driven search (Hyndman & Khandakar, 2008).
@@ 3.2 Global Models
While not being the first ever RNNs, Long Short-term Memory models (LSTMs) (Hochreiter & Schmidhuber, 1997) popularized the concept of iteratively consuming an input sequence to produce both an output and a hidden state propagated to the next step. Using several gating functions, very deep LSTM networks can be trained via backpropagation through time. The recurrent models used in this work are equivalent to the DeepAR model of Salinas et al. (2020). Gated recurrent units (GRUs) simplify LSTMs, often reaching similar performance with fewer parameters (Cho et al., 2014).
@@
It is often beneficial to decompose challenging tasks into a sequence of easier ones. N-BEATS, therefore, comprises a stack of doubly residually connected blocks, where each block performs both a forecast and a backcast (Oreshkin et al., 2020). The backcast is subtracted from the input and only then passed to the subsequent block, in effect removing the component used by the block and only leaving the following block with the residual. Similarly, the complete forecast is computed as the sum of individual forecasts, effectively modeling it as an additive decomposition. Each back- and forecast is computed based on predicting basis function coefficients to ease training and allow for inductive biases. We use the multivariate extension of Darts (Herzen et al., 2022).
@@
xLSTM-Mixer
Recently, mixing emerged as a new paradigm for deep learning architectures, where different dimensions of the data are projected with weight sharing in an alternating fashion (Tolstikhin et al., 2021). This has since been successfully transferred to time series: xLSTM-Mixer (Kraus et al., 2024a) combines recurrent xLSTM modules (Beck et al., 2024) with three different mixing stages to perform forecasting. Namely, the time dimensions are first projected per variate, then time and variate data are jointly projected in the recurrent step, and a final view mixing reconciles a forward and backward view of the series into the output forecast.
Chronos
In language modeling, self-supervised pretraining on vast amounts of data has proven tremendously effective at learning the basic structure of texts (Zhao et al., 2024). This paradigm has recently been transferred to deep time series forecasting (Kraus et al., 2024b). Chronos is one such model closely following the structure of typical transformer-based language models (Ansari et al., 2024). The method is trained on large amounts of real and synthetic time series normalized by mean scaling and discretized like text tokens. The model is both pretrained and deployed by autoregressive discrete next-token prediction. We use the improved variants Bolt Small and Bolt Base.
@@ 4.1.1 Selected Indicators
We experimented with providing the model with additional static context for each company. Namely, we added region information (Americas/Europe/Japan/Rest of Asia) and sector classifications according to the Global Industry Classification Standard (GICS) by MSCI and S&P. We encoded them as numeric one-hot statics for the global models that can use it (see Table 1). However, we only retained the latter since the region information did not noticeably change the forecasts. This is likely due to it being very coarse and of limited relevance in globally diversified companies. See Section A.5 for details.
@@ 4.1.2 Data Preparation
The primary quarterly company fundamentals data was obtained by integrating several proprietary data sources from S&P Global. Considerable manual preprocessing was required to bring information from balance sheets, income statements, and cash flow statements into a single format. In particular, restatements and pro forma/hypothetical statements introduced significant inconsistencies and jumps in the data if not accounted for appropriately. All values were converted into Euros for the sake of comparability. Companies were included if (1) they were publicly traded throughout the entire period and, therefore, published reports, and (2) they had at least 1 billion Euros in market capitalization at some point in the time span. Smaller businesses were excluded because they typically show different dynamics than larger ones. Additional highly atypical companies were removed if they contained more than one data point in any feature that was at least 30 standard deviations from the mean after the normalization described below. We found two rounds of such cleaning of extreme outliers in the long tails (cf. Section A.2) to be sufficient, which removed only 19 companies. This left 2527 companies in the period from 2009 Q1 to 2023 Q3.
Appropriate data representation and normalization are critical in machine learning applications. In the particular case of CFs, one has to counteract the different orders of magnitude of the indicators to ensure more similar numerical ranges of the data. We, therefore, propose an effective domain-specific normalization. For each company individually, all variates derived from the income statement are normalized by Total Revenues, and all variates from the balance sheet are normalized by Total Assets. Total Revenues and Total Assets are divided by each other, respectively. This proved to be more effective than a generic Yeo–Johnson power transformation (Yeo & Johnson, 2000). Finally, the data was normalized to have zero mean and unit variance per feature over all time steps and companies. Details on the resulting challenging dataset, including its moments, high degree of non-stationarity, and diverse seasonality, are provided in Section A.2.
@@ 4.1.3 Results
Figure 1: Global and deep models generally provide better deterministic forecasts than local and pretrained models measured by MSE, but are much more complex. Lower is better. It shows the mean and standard deviation over all five target features, forecast horizon steps, and folds.
@@
The comparison of the models is shown visually in Figure 1. In both metrics, the models fare notably differently. Overall, global models tend to give better results in terms of MSE, while the picture is more varied if performance is measured in sMAPE. Specifically, the univariate local models $\operatorname{ARMA}(4,4)$ and Prophet fail to provide consistent forecasts, as their high MSE and considerable variance hint at some major mispredictions. The pretrained Chronos models fail even more drastically, possibly due to the unusually short length of the time series. While in terms of MSE, the two deep learning RNN models (LSTM and GRU) give the best results, the simplistic local model $\operatorname{ARMean}(1)$— somewhat surprisingly—is best in terms of sMAPE. This highlights the need for forecasts that are not only accurate but also reliable. Therefore, we will now go over to probabilistic forecasts.
Table 2: Recurrent models provide the overall best performance. Comparison of average performance and standard deviation of the models measured by different metrics. The average and standard deviation is taken over all five target features, forecast horizon steps, and folds. The best is marked as bold and the runner-ups as underlined. See also Figure 2 for nCRPS/MAE.
@@
xLSTM-Mixer 0.208±0.63 0.442±22.71 0.233±0.62 3.032±277 8167.136 0.422±0.43 -29758.268 0.146±0.50
Chronos Small 0.225±0.84 0.751±63 0.264±0.83 3.422±344 8673.922 0.421±0.44 -30063.893 0.418±0.76
Chronos Base 0.231±0.86 0.798±57 0.273±0.85 3.403±288 9316.606 0.423±0.44 -30118.321 0.418±0.76
@@
To assess the reliability of the forecasters, we evaluated their ability to quantify uncertainty correctly. Not all models can estimate uncertainty, as shown in Table 1. Sampling forecasts can be achieved in different ways. For instance, one can introduce noise in the state space as in the $\operatorname{ARMA}$, $\operatorname{ARIMA}$, and AutoARIMA models or in the trend component of the Prophet model, which is similar to Monte Carlo dropout in deep learning models (Gal & Ghahramani, 2016). The remaining models can directly estimate the parameters of probability distributions. Therefore, and for a fair comparison, all remaining models perform quantile regression, where the boundaries of certain quantiles were regressed on. Note that these models are trained separately from the ones in the deterministic setting, even though represented jointly in Table 2. We have drawn 100 samples from each probabilistic model and evaluated the resulting empirical distribution against the known test data using the continuous ranked probability score (CRPS) (Gneiting & Raftery, 2007). Intuitively, we first define a helper cumulative distribution function $\mathbf{1}_{\{x\geq y\}}$ over $x$ that steps from zero to one at the scalar value $y$ that we consider the ground truth. The CRPS then measures the integrated squared difference between this helper and the true distribution $E(x)$. It is defined as
@@
Figure 2: Models based on deep learning are clearly superior to all other families. The error of different models when forecasting the normalized features probabilistically. Lower is better. It shows the mean and standard deviation over all five target features, forecast horizon steps, and folds. Since the nCRPS score reduces to the absolute error (MAE) in the case of deterministic forecasts (i.e., a single sample), we include such models for direct comparison marked with . Full results are in Table 2.
An example forecast transformed back to the original feature domain is provided in Figure 3. The normalized forecasts were inverse-transformed back into the original domain using the last-known training data to avoid any information leakage. The model successfully captures the overall dynamics of the variates while tending to lag slightly behind. In particular, the variance of the model appropriately captures the variability of the ground-truth time series.
@@
As shown in Figure 4, not all features are equally predictable by all models. For instance, the Total Equity is easily forecasted by $\operatorname{ARMean}(1)$, yet much harder for others. This is plausible, given that it only moderately changes over time and thus might cause more overfitting. On the other hand, it is surprising that Total Revenues is relatively straightforward to predict for the DL models, indicating that some other covariates are predictive of it. The fact that the Operating Income is more challenging suggests that the cost of revenues varies significantly, affecting the variate in unpredictable ways. Consequently, the Net Income is affected, albeit reduced, possibly due to steadier taxation and interests. By cumulating it over time, one derives the, therefore, equally predictable Cash from Operations.
Figure 4: Models differ substantially in the features they can successfully forecast. This plot shows the mean nCRPS/MAE per model and feature over all folds. Lower is better. Linear Regression and Chronos, with their large errors, are excluded to make differences more discernible. The Operating Income and the Total Equity pose the greatest challenges to the models under investigation.
Global Black Swan events like the COVID-19 pandemic starkly impact forecasting performance, as shown in Figure 5. Note that for most variates and models, this extends far beyond the difficulties in forecasting in the respective periods, like here, 2020 Q1 and onward. On the one hand, starting with 2019 Q2, the forecast evaluation is already infected with COVID-19 effects since we forecast one year into the future. Furthermore, the abnormal data from the pandemic period is then used to learn forecasts for later times, which no longer share the same pandemic dynamics, extending the effect beyond the end of the special period. We can also see that models excel in different parts of the time axis. For example, while $\operatorname{ARMean}(1)$ appeared to be overall decent in Figure 2, it produces terrible forecasts in the more recent time spectrum when we consider uncertainty estimation for Net Income while maintaining excellent performance on Total Equity. There seems to be a visibly recurring pattern in predictability in Cash from Operations every four quarters, where the last quarter of each year is the least predictable.
@@
Figure 6: Forecast error generally rises with increasing look-ahead horizon, with a noticeable drop at four steps. The error is over all features and folds, where lower is better. The shaded area indicates one standard deviation. See Figure 5 for the legend.
@@
We also compared the machine forecasts with those of human expert analysts. To this end, we used data from StarMine by Refinitiv, which contains one-year-ahead non-probabilistic forecasts of the expected Total Revenues for 87.96% of the companies from 2013 Q3 to 2022 Q2. Thus, we need to compare them only with the automatic forecasts for that feature, those companies, the last of the four look-ahead steps, and the valid time window. The results are provided in complete form in Section A.6. In summary, automatic forecasts are superior for most metrics, even though different models than in Table 2 are leading. However, human analysts make mispredictions that better match the scale of the data, e.g., minor errors on small scales and large ones on bigger ground-truth numbers. This results in much better MAPE and sMAPE scores and might hint at potentially more calibrated uncertainty estimates if they were available. Given that human forecasts are commonly employed in investing (Guerard et al., 2015), it is interesting to see how their utility compares to automatic forecasts.